Cilostazol is an antiplatelet anticoagulant. A patient is taking a maintenance dose of 100 mg, which is administered twice a day. Calculate the number of 50 mg tablets required to provide the patient with medication for a week.
Answer #1

Ans) Cilostazol is an antiplatelet anticoagulant. A patient is taking a maintenance dose of 100 mg, which is administered twice a day. Calculate the number of 50 mg tablets required to provide the patient with medication for a week.

100 mg two times in a day means 200 mg

200 × 7 = 1400 mg/ week

50 mg = 1 tab

1400 mg = x

x = 1400/50 = 28 tabs/week.